### v2

@@ Abstract
We consider the problem of designing auctions which maximize consumer surplus (i.e., the social welfare minus the payments charged to the buyers). In the consumer surplus maximization problem, a seller with a set of goods faces a set of strategic buyers with private values, each of whom aims to maximize their own individual utility. The seller, in contrast, aims to allocate the goods in a way which maximizes the total buyer utility. The seller must then elicit the values of the buyers in order to decide what goods to award each buyer. The canonical approach in mechanism design to ensure truthful reporting of the private information is to find appropriate prices to charge each buyer in order to align their objective with the objective of the seller. Indeed, there are many celebrated results to this end when the seller’s objective is welfare maximization [vickrey1961, clarke1971, groves1973] or revenue maximization [myerson1981optimal]. However, in the case of consumer surplus maximization the picture is less clear – using high payments to ensure the highest value bidders are served necessarily decreases their surplus utility, but using low payments may lead the seller into serving lower value bidders.
Our main result in this paper is a framework for designing mechanisms which maximize consumer surplus. We instantiate our framework in a variety of canonical multi-parameter auction settings (i.e., unit-demand bidders with heterogeneous items, multi-unit auctions, and auctions with divisible goods) and use it to design auctions achieving consumer surplus with optimal approximation guarantees against the total social welfare. Along the way, we answer an open question posed by hartline2008optimal, who, to our knowledge, were the first to study the question of consumer surplus approximation guarantees in single-parameter settings, regarding optimal mechanisms for two bidders.
@@ 1 Introduction
The use of monetary transfers is arguably the most ubiquitous approach in mechanism design for incentivizing truthful reporting of private information. In our single-item auction setting, for example, the well-known second-price auction, which awards the item to the bidder with the highest reported value and charges this bidder a price equal to the second highest reported value, achieves exactly the goal of our auctioneer. Namely, all bidders are incentivized to report their true values to the auctioneer and the social welfare is maximized. By using transfers of money from the agents to the designer, one is able to design mechanisms with similar properties for a broad class of allocation problems. Indeed, when an auctioneer aims to maximize social welfare the celebrated Vickrey-Clarke-Groves (VCG) mechanism [vickrey1961, clarke1971, groves1973] works in totally general allocation environments, where an auctioneer aims to allocate a set of goods and the bidders can have arbitrary valuations over the subsets of items. By carefully constructing appropriately large payments from the bidders to the auctioneer, the VCG mechanism ensures that bidders truthfully report all their private information and exactly maximizes the social welfare.
@@
On the other hand, there are many scenarios in which a mechanism designer may wish to limit or reduce the payments she collects from the bidders. First, in some settings monetary payments may be undesirable or even illegal and, as such, alternative forms of “payment” which offer no real revenue to the auctioneer are employed. For instance, so-called “ordeal mechanisms” for allocating welfare benefits [alatas2013ordeal, nichols1982targeting, zeckhauser2021strategic, sylvia2022ordeal] may utilize long applications or waiting time in a queue. In these mechanisms time and effort stand in for money as a form of payment. Second, in “proof of work” environments [dwork1992pricing, jakobsson1999proofs], such as the Bitcoin blockchain, agents all want access to some shared resource, e.g., the rights to publish the next block, and effort in finding a solution to a difficult computational problem is used to decide which agent(s) are awarded access. This solution to the computational problem has no value and effort is merely used as “payment” to coordinate the agents. Third, in many allocation settings involving computing resources the “payments” come in the form of a reduced quality of service. For instance, packets may be delayed by servers [cole2003much] or some of the resources may go unused [cole2013mechanism].
@@
In each of the aforementioned settings the payments have no direct benefits outside of ensuring good incentives and, as such, the payments from the bidders are effectively “burned”. Therefore, a natural goal is to find mechanisms that: (i) still ensure truthful reporting, (ii) find outcomes of high welfare, and (iii) keeps the total payments low. Notably, however, although the VCG mechanism is truthful and maximizes the social welfare, satisfying the first two goals, the payments it specifies can be quite high and may leave, essentially, no residual consumer surplus utility for the bidders themselves. For instance, in the second-price auction when the second highest value is nearly the same as the first highest, the consumer surplus is nearly 00.
@@
Aiming to address this issue, hartline2008optimal introduced the study of “money-burning mechanisms” – mechanisms which may charge payments from the bidders but seek to maximize the residual consumer surplus. Focusing on the single-item setting, they identify optimal Bayesian mechanisms (i.e., optimal mechanisms when bidders draw their values from distributions known to the auctioneer) for consumer surplus. They then extend these results into worst-case prior-free settings by giving a truthful mechanism which always obtains consumer surplus within a $O(\log(n))$-factor of the total social welfare and demonstrate that this is the best possible. Finally, they expand their results from single-item settings to single-parameter $k$-unit settings where an auctioneer has $k$ identical items and each bidder aims to receive one of these $k$ items.
@@
From the groundbreaking work of hartline2008optimal, many natural questions arise. In particular, they call out two specific questions, namely, (i) “quantify[ing] the power of money burning” in “settings beyond i.i.d. distributions and $k$-unit [unit-demand] auctions”; and (ii) finding improved “upper and lower bounds for … money-burning mechanisms with a small number of agents”, where the “best-possible approximation ratio … even in the two agent, one unit special case” remained unknown. In this work, we tackle both these questions and (partially) resolve them. We study three well-established multi-parameter auction settings which all generalize the single-item setting studied in [hartline2008optimal] and give mechanisms achieving asymptotically optimal consumer surplus guarantees. We then narrow our attention to the two-bidder single-item setting and identify exactly optimal mechanisms therein.
Our Contribution and Techniques
Unlike in the case of welfare maximization, where we can use the VCG mechanism directly, the difficulty in maximizing residual consumer surplus involves carefully balancing the prices. When prices are set too high, all bidders will have low utility. On the other hand, if prices are set too low, the auctioneer may serve bidders of low total welfare. To overcome this problem we introduce a framework for designing mechanisms for consumer surplus maximization which we call the “VCG with copies mechanism” (Section 2.2). In this framework, we essentially run the VCG mechanism over a modified instance instead of the original one. In the modified instance, we have multiple copies of each item, and the agents’ valuations are modified so that each agent will not (strictly) prefer a set of items that were not available in the original instance. This, essentially, results in more balanced prices which are lower than the VCG prices for the original instance.
@@
The framework consists of three parts. The first is deciding on the number of copies, or more generally choosing a distribution over the number of copies. Then, the VCG mechanism is run over the modified instances with the chosen number of copies. The last part is a rounding scheme that takes the VCG allocation and payments for the modified instance and turns them into an allocation and payments for the original instance where we only have a single copy of each item. Implementing this last step can be tricky as we want to convert the allocation and payments such that they are still relatively close to the allocation and payments of the modified instance while maintaining truthfulness.
We instantiate this framework with three multi-parameter auction settings, and obtain optimal approximation guarantees against the social welfare in each setting.
@@
First, we consider multi-unit auctions with bidders whose valuation functions are submodular. In this case, using the VCG with copies framework we get a mechanism whose residual surplus is a $O(\log n)$ approximation to the social welfare, where $n$ is the number of bidders (Theorem 8).
Second, we consider an auction setting with heterogeneous items and unit-demand buyers. In the prior free case, using the framework, we get a mechanism whose residual surplus is a $O(\log n)$ approximation to the social welfare (Theorem 7). We also consider a Bayesian case, where for every agent, his value for each item is sampled $i.i.d$. In this case, using the framework, we get a better approximation guarantee of $O(1+\log\frac{n}{m})$ (Theorem 9), which generalizes the result of [hartline2008optimal] for $m$-unit auction with $n$ unit-demand agents of $O(1+\log\frac{n}{m})$ approximation. In fact, we need a weaker requirement than the assumption that for every agent his value for each item is sampled $i.i.d$. We use only the assumption that the item with the highest value of each agent will be distributed independently uniformly at random, which we call the uniform highest value assumption.
hartline2008optimal showed a lower bound of $\Omega(1+\log\frac{n}{m})$ for multi-unit auctions with unit demand bidders. Building on their lower bound instance, we construct two instances. In the first we have $n$ additive bidders, each bidder’s value for the grand bundle is sampled independently from the exponential distribution with parameter $1$, and his value for a fraction of the grand bundle is a fraction of this value. In the second instance, all bidders are only interested in the first item, and their value for this item is sampled independently from the exponential distribution with parameter $1$. Very similar to the proof of [hartline2008optimal], we get a lower bound of $\Omega(\log n)$ for both instances. The first instance shows the tightness of our result for multi-unit auctions with submodular bidders, and the second shows tightness for the case of prior free unit demand agents.
@@
In this paper, we also study the special case of two agents (Section 4). First, we show that the case of a single-item is as hard as any number of items when agents have arbitrary monotone valuations; in both cases an approximation of $\frac{3}{2}$ is optimal. This result is consistent with our previously discussed results which show that the surplus guarantees achievable in several well-studied multi-parameter auction settings with many agents are no worse than the surplus guarantees achievable in the canonical single-item setting with many agents. Then, in Section 4.1, we answer an open question that was raised by hartline2008optimal, regarding the approximation ratio with respect to a different benchmark that is not the unattainable first-best.
@@ 1.1 Additional Related Work
In addition to [hartline2008optimal], the works perhaps closest to ours are [fotakis2016efficient, goldner2024simple]. fotakis2016efficient study maximizing consumer surplus in fully general domains, but obtain only a $O(\log|\mathcal{O}|)$ approximation guarantee where $\mathcal{O}$ is the set of feasible outcomes. Since $|\mathcal{O}|$ can be exponential in the number of items in our indivisible goods settings (and, essentially infinite in our divisible good settings), while their mechanisms apply in our settings, they guarantee only a trivial fraction of the social welfare as consumer surplus (which can be achieved simply by following [hartline2008optimal] on the grand bundle of items). Seeking to improve upon the guarantees (in terms of bidders and items) of [fotakis2016efficient], goldner2024simple initiated the study of sublinear approximations in multi-dimensional settings. Independently and concurrently to our work, they studied multi-dimensional surplus maximization with unit-demand bidders in Bayesian setting as we do in Section 2.4. For the case where bidders’ values are distributed i.i.d., they devise a Bayesian incentive compatible (rather than truthfulness in expectation) auction. In Section 2.4, we show that our mechanism achieves the same asymptotic guarantee as [goldner2024simple] under a weaker assumption where for each agent the favorite item is distributed uniformly among the items (which generalizes both the i.i.d. case considered in [goldner2024simple] and the identical items unit-demand case considered in [hartline2008optimal]).
@@
Also quite similar to our work is that of [qiao2023online, ganesh2023combinatorial]. qiao2023online propose the problem of “pen-testing”. In the pen-testing problem, an algorithm faces a set of $n$ pens each with an unknown amount of ink. The goal of the algorithm is to “test” these pens, possibly exhausting some of the ink contained within, in order to select a pen with a large amount of remaining ink. As qiao2023online and ganesh2023combinatorial argue, one can interpret this problem as a consumer surplus maximization problem where the pens are the bidders. qiao2023online then give algorithms for selecting a single pen in online pen-testing settings (where pens arrive one-by-one and the algorithm must make irrevocable decisions). ganesh2023combinatorial consider the problem of selecting multiple pens in both online and offline environments where a feasibility constraint restricts which subsets of pens can be simultaneously selected. Notably, these works only apply in binary, single-parameter mechanism design settings, since choosing or not choosing a pen corresponds to a binary service decision for the bidders (i.e., each bidder has a single private value for receiving a service and either receives the service or not), whereas the problems we investigate are all multi-parameter auction settings (i.e., bidders hold multiple pieces of private information).
@@
Similar in spirit to our question of mechanism design with money-burning is the literature on mechanism design without money. Mechanism design without money is a rich literature (see [schummer2007mechanism] for a survey of some results) seeking to design mechanisms which align the incentives of the agents with the designer while avoiding monetary payments altogether. Indeed, one possible approach to the questions we explore in this work is to find a welfare maximizing mechanism without money. However, it is known from prior work, e.g., [hartline2008optimal], that mechanisms without money cannot guarantee better than an $\Omega(n)$ approximation to the optimal social welfare, even in the case of allocating a single good. As such, we seek mechanisms which leverage (small) monetary transfers in order to obtain better consumer surplus guarantees.
@@
Finally, money burning sees practical application in the use of transaction fee mechanisms in blockchain (see, e.g., [roughgarden2021transaction, chung2023foundations, chen2023bayesian, wu2023maximizing]). Since in a blockchain protocol only a limited number of transactions can be posted to the chain per block, transaction fee mechanisms are used to decide which transactions are ultimately processed. In these mechanisms some of the payments are burnt and are removed from the system altogether, rather than transferred to the miner who is processing the block. In contrast to our problem, however, payments are burnt in transaction fee mechanisms in order to incentivize the miners (who are also assumed to be strategic agents in the model) to follow the blockchain protocol. As such, the literature on transaction fee mechanism design centers, generally, only around designing mechanisms with good incentive properties (and can allow payments to be quite high), whereas in this work we aim to satisfy a simpler set of incentives while optimizing the more challenging objective of consumer surplus.

### v3

@@ Abstract
We consider the problem of designing auctions that maximize consumer surplus (i.e., the social welfare minus the payments charged to the buyers). In the consumer surplus maximization problem, a seller with a set of goods faces a set of strategic buyers with private values, each of whom aims to maximize their own individual utility. The seller, in contrast, aims to allocate the goods in a way that maximizes the total buyer utility. The seller must then elicit the values of the buyers in order to decide what goods to award each buyer. The canonical approach in mechanism design to ensure truthful reporting of the private information is to find appropriate prices to charge each buyer in order to align their objective with the objective of the seller. Indeed, there are many celebrated results to this end when the seller’s objective is welfare maximization [vickrey1961, clarke1971, groves1973] or revenue maximization [myerson1981optimal]. However, in the case of consumer surplus maximization the picture is less clear – using high payments to ensure the highest value bidders are served necessarily decreases their surplus utility, but using low payments may lead the seller into serving lower value bidders.
Our main result in this paper is a framework for designing mechanisms that maximize consumer surplus. We instantiate our framework in a variety of canonical multi-parameter auction settings (i.e., unit-demand bidders with heterogeneous items, multi-unit auctions, and auctions with divisible goods) and use it to design auctions achieving consumer surplus with tight approximation guarantees against the total social welfare. Along the way, we resolve an open question posed by hartline2008optimal for the two bidder single item setting.
@@ 1 Introduction
The use of monetary transfers is arguably the most ubiquitous approach in mechanism design for incentivizing truthful reporting of private information. In our single-item auction setting, for example, the well-known second-price auction, which awards the item to the bidder with the highest reported value and charges this bidder a price equal to the second highest reported value, achieves exactly the goal of our auctioneer. Namely, all bidders are incentivized to report their true values to the auctioneer and the social welfare is maximized. By using transfers of money from the agents to the designer, one is able to design mechanisms with similar properties for a broad class of allocation problems. Indeed, when an auctioneer aims to maximize social welfare the celebrated Vickrey-Clarke-Groves (VCG) mechanism [vickrey1961, clarke1971, groves1973] works in settings with multiple items even if the bidders have arbitrarily complex valuations over subsets of items. By carefully constructing appropriately large payments from the bidders to the auctioneer, the VCG mechanism ensures that bidders truthfully report all their private information and exactly maximizes the social welfare.
@@
On the other hand, there are many scenarios in which a mechanism designer may wish to limit or reduce the payments she collects from the bidders. First, in some settings monetary payments may be undesirable or even illegal and, as such, alternative forms of “payment” which offer no real revenue to the auctioneer are employed. For instance, so-called “ordeal mechanisms” for allocating welfare benefits [alatas2013ordeal, nichols1982targeting, zeckhauser2021strategic, sylvia2022ordeal] may utilize long applications or waiting time in a queue. In these mechanisms time and effort stand in for money as a form of payment. Second, in “proof of work” environments [dwork1992pricing, jakobsson1999proofs], such as the Bitcoin blockchain, agents all want access to some shared resource, e.g., the rights to publish the next block, and effort in finding a solution to a difficult computational problem is used to decide which agent(s) are awarded access. This solution to the computational problem has no value and effort is merely used as “payment” to coordinate the agents. Third, in many allocation settings with computing resources the “payments” come in the form of reduced service quality. For instance, packets may be delayed by servers [cole2003much] or some resources may go unused [cole2013mechanism].
@@
In each of the aforementioned settings the payments have no direct benefits outside of ensuring good incentives and, as such, the payments from the bidders are effectively “burned”. Therefore, a natural goal is to find mechanisms that: (i) still ensure truthful reporting, (ii) find outcomes of high welfare, and (iii) keep the total payments low. Notably, however, although the VCG mechanism is truthful and maximizes the social welfare, satisfying the first two goals, the payments it specifies can be quite high and may leave, essentially, no residual consumer surplus utility for the bidders themselves. For instance, in the second-price auction when the second highest value is nearly the same as the first highest, the consumer surplus is nearly 00.
@@
Aiming to address this issue, hartline2008optimal introduced the study of “money-burning mechanisms” – mechanisms which may charge payments from the bidders but seek to maximize the residual consumer surplus. Focusing on the single-item setting, they identify optimal Bayesian mechanisms (i.e., optimal mechanisms when bidders draw their values from distributions known to the auctioneer) for consumer surplus. Then, they extend these results into worst-case prior-free settings by giving a truthful mechanism that always obtains consumer surplus within a $O(\log(n))$-factor of the total social welfare where $n$ is the number of bidders and demonstrate that this is the best possible. Finally, they expand their results from single-item settings to single-parameter $m$-unit settings where an auctioneer has $m$ identical items and each bidder aims to receive one of these $m$ items.
@@
From the groundbreaking work of hartline2008optimal, many natural questions arise. In particular, they call out two specific questions, namely, (i) “quantify[ing] the power of money burning” in “settings beyond i.i.d. distributions and $k$-unit [unit-demand] auctions”; and (ii) finding improved “upper and lower bounds for … money-burning mechanisms with a small number of agents”, where the “best-possible approximation ratio … even in the two agent, one unit special case” remained unknown. In this work, we tackle both these questions and (partially) resolve them. We study three well-established multi-parameter auction settings which all generalize the single-item setting studied in [hartline2008optimal] and give mechanisms achieving asymptotically tight consumer surplus guarantees. We then narrow our attention to the two-bidder single-item setting and identify exactly the optimal mechanisms therein.
1.1 Our Contribution and Techniques
Maximizing residual consumer surplus poses challenges that do not arise in welfare maximization, where the VCG mechanism can be applied directly. The difficulty lies in balancing prices: setting them too high reduces bidders’ utility, while setting them too low risks allocating items inefficiently.
To address this challenge, we introduce the VCG with copies framework (Section 2.2). The key idea behind this framework is to run the VCG mechanism on a modified instance instead of the original one. In this modified instance, multiple copies of each item are introduced, and agents’ valuations are adjusted so that they do not strictly prefer allocations that were unavailable in the original setting. This approach effectively results in more balanced prices that are lower than those produced by VCG in the original instance.
@@ 1.1 Our Contribution and Techniques
The framework consists of four steps. First, the number of copies, or a distribution over possible numbers of copies, is determined. Second, the VCG mechanism is run over the modified instance. Third, a rounding scheme takes the VCG allocation for the modified instance and turns it into an allocation for the original instance where we only have a single copy of each item. The final step adjusts the VCG payments proportionally to the rounded allocation. The third step of using a rounding scheme is nontrivial, as it must ensure that the final allocation remains close to the allocation in the modified instance while maintaining truthfulness.
We instantiate this framework across three multi-parameter auction settings, achieving tight approximation guarantees against the social welfare  [^1] All our lower bounds follow from the bound in [hartline2008optimal].. The resulting mechanisms are prior-free, while the matching lower bounds hold even for Bayesian mechanisms [^2] These lower bounds hold for instances where agent values are independently drawn from a common distribution, and mechanisms need only be Bayesian incentive compatible..
@@
First, we analyze multi-unit auctions where bidders have submodular valuation functions. We show that applying our framework in this setting yields a tight logarithmic approximation to the optimal welfare (Theorem 8).
Next, we consider unit-demand bidders. For adversarial heterogeneous items, we show that an $O(\log n)$ approximation to the welfare is achievable (Theorem 7), and this bound is tight due to [hartline2008optimal]. In Bayesian settings, we refine our analysis of the mechanism’s approximation guarantee. Specifically, if the distributions of the agents’ valuations satisfy the no-superstar-item assumption that assumes the probability of any item being the favorite item of some agent is at most $\frac{c}{m}$, we obtain an improved approximation of $O(\log(1+\frac{n\cdot c}{m}))$ (Theorem 9). This generalizes the results of both [hartline2008optimal] and [goldner2024simple]. In particular, hartline2008optimal showed that for identical items there exists a mechanism whose residual surplus provides an $O\left(\log\left(1+\frac{n}{m}\right)\right)$ approximation to the welfare, and this bound is asymptotically tight. Later, goldner2024simple considered the i.i.d. case, where they presented a simple mechanism obtaining the same approximation (of the welfare) of $O\left(\log\left(1+\frac{n}{m}\right)\right)$, which is tight. Notably, both are special cases of our no-superstar-item assumption with $c=1$. At the other extreme, when $c=m$ (corresponding to the adversarial heterogeneous items), we recover our bound for heterogeneous items. To devise our mechanism, we take a different approach from [hartline2008optimal] and [goldner2024simple]. hartline2008optimal utilize the single-parameter nature of the problem, while the mechanism of goldner2024simple relies heavily on the symmetry induced by the i.i.d. assumption, where the truthfulness is not maintained once the i.i.d. assumption is relaxed. Figure 1 summarizes this hierarchy of the complexity of different variants of the problem.
[…]
Figure 1: Hierarchy of the complexity of different variants for unit-demand bidders.
@@
In this paper, we also study the special case of two agents (Section B). First, we show that the case of a single-item is as hard as any number of items when agents have arbitrary monotone valuations; in both cases an approximation of $\frac{3}{2}$ is optimal. This result is consistent with our previously discussed results which show that the surplus guarantees achievable in several well-studied multi-parameter auction settings with many agents are no worse than the surplus guarantees achievable in the canonical single-item setting with many agents. Then, in Section B.1, we answer an open question that was raised by hartline2008optimal, regarding the approximation ratio with respect to a different benchmark that is not the unattainable first-best.
@@ 1.2 Additional Related Work
In addition to [hartline2008optimal], the works perhaps closest to ours are [fotakis2016efficient, goldner2024simple]. fotakis2016efficient study maximizing consumer surplus in fully general domains, but obtain only a $O(\log|\mathcal{O}|)$ approximation guarantee where $\mathcal{O}$ is the set of feasible outcomes. Since $|\mathcal{O}|$ can be exponential in the number of items in our indivisible goods settings (and, essentially infinite in our divisible good settings), while their mechanisms apply in our settings, they guarantee only a trivial fraction of the social welfare as consumer surplus (which can be achieved simply by following [hartline2008optimal] on the grand bundle of items). Seeking to improve upon the guarantees (in terms of bidders and items) of [fotakis2016efficient], goldner2024simple initiated the study of sublinear approximations in multi-dimensional settings. Independently of our work, they studied multi-dimensional surplus maximization with unit-demand bidders in Bayesian settings as we do in Section 2.4. For the case where bidders’ values are distributed i.i.d., they devise a Bayesian incentive compatible (rather than truthful in expectation) auction. In Section 2.4, we show that our mechanism achieves the same asymptotic guarantee as [goldner2024simple] under a weaker assumption where the probability of each item being the favorite of some agent is bounded by $\frac{c}{m}$ (which generalizes both the i.i.d. case considered in [goldner2024simple] and the identical items unit-demand case considered in [hartline2008optimal] when setting $c=1$).
@@
Also quite similar to our work is that of [qiao2023online, ganesh2023combinatorial]. qiao2023online propose the problem of “pen-testing”. In the pen-testing problem, an algorithm faces a set of $n$ pens each with an unknown amount of ink. The goal of the algorithm is to “test” these pens, possibly exhausting some of the ink contained within, in order to select a pen with a large amount of remaining ink. As qiao2023online and ganesh2023combinatorial argue, one can interpret this problem as a consumer surplus maximization problem where the pens are the bidders. qiao2023online then give algorithms for selecting a single pen in online pen-testing settings (where pens arrive one-by-one and the algorithm must make irrevocable decisions). ganesh2023combinatorial consider the problem of selecting multiple pens in both online and offline environments where a feasibility constraint restricts which subsets of pens can be simultaneously selected. Notably, these works only apply in binary-decision, single-parameter mechanism design settings, since choosing or not choosing a pen corresponds to a binary service decision for the bidders (i.e., each bidder has a single private value for receiving a service and either receives the service or not), whereas the problems we investigate are all multi-parameter auction settings (i.e., bidders hold multiple pieces of private information).
@@
Similar in spirit to our question of mechanism design with money-burning is the literature on mechanism design without money. Mechanism design without money is a rich literature (see [schummer2007mechanism] for a survey of some results) seeking to design mechanisms that align the incentives of the agents with the designer while avoiding monetary payments altogether. Indeed, one possible approach to the questions we explore in this work is to find a welfare maximizing mechanism without money. However, it is known from prior work, e.g., [hartline2008optimal], that mechanisms without money cannot guarantee better than an $\Omega(n)$ approximation to the optimal social welfare, even in the case of allocating a single good. As such, we seek mechanisms that leverage (small) monetary transfers in order to obtain better consumer surplus guarantees.
@@
Another work that resembles our approach considers mechanisms with modified supplies of goods. In particular, roughgarden2012supply design mechanisms that restrict the capacities of goods, whereas our approach takes the opposite route: we deliberately allocate additional goods and then apply rounding techniques to prevent over-allocation. While the former has proven useful for revenue maximization, we demonstrate that the latter is effective for maximizing surplus.
Finally, money burning sees practical application in the use of transaction fee mechanisms in blockchain (see, e.g., [roughgarden2021transaction, chung2023foundations, chen2023bayesian, wu2023maximizing]). Since in a blockchain protocol only a limited number of transactions can be posted to the chain per block, transaction fee mechanisms are used to decide which transactions are ultimately processed. In these mechanisms, some of the payments are burnt and are removed from the system altogether, rather than transferred to the miner who is processing the block. In contrast to our problem, however, payments are burnt in transaction fee mechanisms in order to incentivize the miners (who are also assumed to be strategic agents in the model) to follow the blockchain protocol. As such, the literature on transaction fee mechanism design centers, generally, only around designing mechanisms with good incentive properties (and can allow payments to be quite high), whereas in this work we aim to satisfy a simpler set of incentives while optimizing the more challenging objective of consumer surplus.